# SoloSpeech: Enhancing Intelligibility and Quality in Target Speech Extraction through a Cascaded Generative Pipeline

### v2

@@ H.1 Comparison of audio compressors
Table 7: Reconstruction results of different audio compressors on the Libri2Mix dataset.
Data Model STFT-D $\downarrow$ Mel-D $\downarrow$ PESQ $\uparrow$ ESTOI $\uparrow$ SISNR $\uparrow$
Clean Single Stable Audio VAE [30] $0.540$ $0.231$ $4.151$ $0.984$ $17.094$
T-F Audio VAE (ours) $\boldsymbol{0.432}$ $\boldsymbol{0.116}$ $\boldsymbol{4.506}$ $\boldsymbol{0.995}$ $\boldsymbol{18.379}$
Noisy mixture Stable Audio VAE [30] $0.570$ $0.147$ $4.327$ $0.984$ $14.575$
T-F Audio VAE (ours) $\boldsymbol{0.450}$ $\boldsymbol{0.076}$ $\boldsymbol{4.529}$ $\boldsymbol{0.995}$ $\boldsymbol{15.571}$
H.2 Comparison of different settings of the audio compressor
We conducted experiments to evaluate the reconstruction quality under different audio compressor settings, including variations in frame rate and latent dimension. Table 8 summarizes the impact of these configurations. Lower frame rates and smaller latent dimensions generally make it easier to train latent diffusion models. Ultimately, we select a frame rate of 50 Hz and a latent dimension of 128 as a trade-off between training efficiency and reconstruction quality.
@@ H.2 Comparison of different settings of the audio compressor
Table 8: Impact of different settings for the audio compressor on the noisy mixture.
Frame Rate Latent Dim STFT-D $\downarrow$ Mel-D $\downarrow$ PESQ $\uparrow$ ESTOI $\uparrow$ SISNR $\uparrow$
25 Hz 64 $0.596$ $0.182$ $3.979$ $0.978$ $14.105$
128 $0.522$ $0.133$ $4.204$ $0.988$ $14.996$
50 Hz 64 $0.510$ $0.128$ $4.268$ $0.989$ $15.080$
128 $0.450$ $0.076$ $4.529$ $0.995$ $15.571$
100 Hz 64 $0.454$ $0.075$ $4.510$ $0.995$ $15.499$
128 $0.440$ $0.071$ $4.530$ $0.995$ $15.688$
H.3 Comparison of Masking Ratios
@@
H.4 Comparison of audio durations
SoloSpeech was trained on audio segments of varying lengths, ranging from 3 to 10 seconds. The model also supports longer input thanks to the Rotary Position Embeddings (RoPE) in the Transformer backbone. In our experiments on Libri2Mix, CHiME-5, and RealSEP, the input segments range from 3 to 20 seconds, and we observe no significant degradation in performance as the input length increases, which is shown in Table 10. This indicates that SoloSpeech is robust to variations in audio duration.
Table 9: Comparison of different audio durations.
@@ H.4 Comparison of audio durations
H.5 Comparison of different model sizes of the target extractor
We present results for different model sizes of the target extractor in Table 10.
Table 10: Comparison of different model sizes of the target extractor. No corrector is used.
Model Size SISNR $\uparrow$ DNSMOS $\uparrow$ WER $\downarrow$
Small $7.68$ $3.71$ $0.26$
Base $8.10$ $3.76$ $0.24$
Large $8.40$ $3.77$ $0.22$
Appendix I Visualization analysis
@@
(a) Sample I: Ground truth
(b) Sample I: SoloSpeech
(c) Sample I: USEF-TSE
(d) Sample II: Ground truth
(e) Sample II: SoloSpeech
(f) Sample II: USEF-TSE
(g) Sample III: Ground truth
(h) Sample III: SoloSpeech
(i) Sample III: USEF-TSE
(j) Sample IV: Ground truth
(k) Sample IV: SoloSpeech
(l) Sample IV: USEF-TSE
(m) Sample V: Ground truth
(n) Sample V: SoloSpeech
(o) Sample V: USEF-TSE
Figure 10: Comparison of the spectrograms of the ground truth, audio extracted by SoloSpeech and by USEF-TSE.
To analyze the impact of the corrector, we present three visualization examples in Figure 11. The corrector primarily functions by reducing background noise and artifacts in the high frequencies. Details of these samples are as follows:
  • Sample I: The corrector reduces background noise artifacts and better reconstructs frequency details. Without the corrector, there is an obvious mispronunciation in the middle of the speech.
  • Sample II: The corrector effectively reduces artifacts, especially affecting the last phoneme in this speech.
  • Sample III: Using the corrector results in less noise production and clearer frequency bands.
(a) Sample I: Ground truth
(b) Sample I: w/ corrector
(c) Sample I: w/o corrector
(d) Sample II: Ground truth
(e) Sample II: w/ corrector
(f) Sample II: w/o corrector
(g) Sample III: Ground truth
(h) Sample III: w/ corrector
(i) Sample III: w/o corrector
Figure 11: Comparison of the spectrograms of the ground truth, audio extracted by SoloSpeech with corrector and by SoloSpeech without corrector.

### v3

@@ V-E3 Real-world Results
TABLE IV: Reconstruction results of different audio compressors on the Libri2Mix dataset.
Data Method PESQ ESTOI SISNR
Clean Single Stable Audio VAE [51] $4.151$ $0.984$ $17.094$
T-F Audio VAE (ours) $\boldsymbol{4.506}$ $\boldsymbol{0.995}$ $\boldsymbol{18.379}$
Noisy mixture Stable Audio VAE [51] $4.327$ $0.984$ $14.575$
T-F Audio VAE (ours) $\boldsymbol{4.529}$ $\boldsymbol{0.995}$ $\boldsymbol{15.571}$
V-F Ablation studies
V-F1 Comparison of audio compressors
In the audio compressors reconstruction experiments, we utilized a set of audio quality metrics: PESQ, ESTOI, and SI-SNR. As shown in Table IV, the T-F Audio VAE consistently outperforms the Stable Audio VAE across all metrics, demonstrating superior audio reconstruction quality.
@@ V-F1 Comparison of audio compressors
TABLE V: Impact of different settings for the audio compressor on the noisy mixture.
Frame Rate Latent Dim PESQ ESTOI SISNR
25 Hz 64 $3.979$ $0.978$ $14.105$
128 $4.204$ $0.988$ $14.996$
50 Hz 64 $4.268$ $0.989$ $15.080$
128 $4.529$ $0.995$ $15.571$
100 Hz 64 $4.510$ $0.995$ $15.499$
128 $4.530$ $0.995$ $15.688$
We conducted experiments to evaluate the reconstruction quality under different audio compressor settings, including variations in frame rate and latent dimension. Table V summarizes the impact of these configurations. Lower frame rates and smaller latent dimensions generally make it easier to train latent diffusion models. Ultimately, we select a frame rate of 50 Hz and a latent dimension of 128 as a trade-off between training efficiency and reconstruction quality.
In addition, we compared the proposed T-F Audio VAE with the state-of-the-art Stable Audio VAE. We employed the VAEs as audio compressors within SoloSpeech and assessed their impact on the TSE task. As presented in Table III, the newly proposed T-F Audio VAE significantly enhances TSE performance compared to the Stable Audio VAE.
V-F2 Comparison of speaker conditions
We compared different speaker conditions and reported the results in Table III. Specifically, We evaluated a speaker embedding [^12] https://github.com/TaoRuijie/ECAPA-TDNN-based approach with AdaLN [25], considering both fixed and fine-tuned configurations; an SSL-based speaker network [^13] https://github.com/BUTSpeechFIT/wespeaker_ssl_public [43]; a time-domain speaker network [^14] https://github.com/BUTSpeechFIT/speakerbeam [66]; and our proposed VAE latent-domain speaker network. Our method does not require additional data to train speaker embeddings, yet achieves the best results among all methods. This superiority is attributed to our approach’s ability to preserve the sequential information of the conditional features and prevent potential mismatching between VAE latents and other types of features representing the target speech.
V-F3 Impact of the corrector
We investigated the impact of the corrector and present the results in Table III. Our proposed corrector significantly outperforms the previous work, Fast-GeCo, thereby substantially enhancing perceptual quality and intelligibility. The corrector contributes significantly to SI-SNR and WER gains, as we found that it can significantly improve signal quality and reduce artifacts in high frequencies, which may influence the ASR system. Furthermore, we applied the proposed corrector to two prior baselines, USEF-TSE (a discriminative method) and SoloAudio (a generative method), and reported the results in Table III. The corrector consistently improved performance across all evaluation metrics. Notably, SoloSpeech still outperforms USEF-TSE even with the corrector applied, especially in DNSMOS scores, highlighting the advantage of generative models in producing more natural-sounding speech. Compared to SoloAudio with the same corrector, SoloSpeech achieves significantly better perceptual quality and intelligibility, demonstrating the effectiveness of our target extractor backbone.
V-F4 Comparison of Masking ratios
@@
V-F5 Comparison of audio durations
SoloSpeech was trained on audio segments of varying lengths, ranging from 3 to 10 seconds. The model also supports longer input thanks to the Rotary Position Embeddings (RoPE) in the Transformer backbone. In our experiments on Libri2Mix, CHiME-5, and RealSEP, the input segments range from 3 to 20 seconds, and we observe no significant degradation in performance as the input length increases, which is shown in Table VI. This indicates that SoloSpeech is robust to variations in audio duration.
TABLE VI: Comparison of different audio durations.
@@ V-F5 Comparison of audio durations
VI Visualization analysis
@@
Figure 6: Comparison of the spectrograms of the ground truth, audio extracted by SoloSpeech, and by USEF-TSE. Rows: Sample I–V. Columns: (a,d,g,j,m) Ground truth, (b,e,h,k,n) audio extracted by SoloSpeech, (c,f,i,l,o) audio extracted by USEF-TSE.
VII Conclusion and Discussions
We proposed a cascaded generative pipeline that offers a strong foundation for future research in generative TSE. Each component is validated for its effectiveness and designed with modularity and scalability in mind.
Compared to previous state-of-the-art methods in both TSE and SS tasks, SoloSpeech achieves significantly higher perceptual quality, naturalness, and intelligibility. Furthermore, when evaluated on out-of-domain data, SoloSpeech demonstrates excellent generalization and robustness.
In real-world evaluations, we observe that strong reverberation and moving sound sources, such as those in the RealSEP dataset, remain challenging, and we leave these issues to future work. We also plan to explore more efficient backbone architectures for each component.